OpenStudy (anonymous):

solve the following system of equations: use the substitution method{5x+6y+8=0 {3x-2y+16=0

OpenStudy (anonymous):

Re-write equation 1, solving for y 5x+8=-6y 6y= -8 - 5x y= -4/3 - 5x/6 Now substitute this value for y into the second equation so that 3x-2y+16=0, becomes 3x-(-4/3 - 5x/6) +16 = 0, then expand the negative sign into the brackets 3x +4/3 +5x/6 +16 = 0, Simplify 23x/6 + 52/3 = 0 23x/6 = -52/3 23x/2 = -52 23x = -104 x = -104/23

OpenStudy (anonymous):

Then substitute that value for x back into the first equation to find y y=-4/3 - x/6